### v1

# Evaluation of Audio Compression Codecs

abstract=true, fontsize=10pt, footinclude=false, headinclude=false, headsepline=false, mpinclude=false, numbers=noenddot, paper=letter, parskip=half-, twocolumn=true, twoside=false, [subfloat]aboveskip=0pt, belowskip=0pt, font=scriptsize,sf, format=hang, labelsep=space, margin=1pt, singlelinecheck=0,
Evaluation of Audio Compression Codecs
Thien T. Duong    Jan P. Springer
Abstract
Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity. It is how humans respond to the accuracy, intelligibility, and fidelity of aural media. Today this fidelity is also heavily influenced by the use of audio compression codecs for storing aural media in digital form. We argue that, when choosing an audio compression codec, users should not only look at compression efficiency but also consider the sonic perceptual quality properties of available audio compression codecs.
We evaluate several commonly used audio compression codecs in terms of compression performance as well as their sonic perceptual quality via codec performance measurements, visualizations, and PEAQ scores. We demonstrate how perceptual quality is affected by digital audio compression techniques, providing insights for users in the process of choosing a digital audio compression scheme.
Keywords: Audio Codecs, Audio Compression, Information Visualization, Quantitative and Qualitative Audio Evaluation, PEAQ
1 Introduction
Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity [Ellis:PRA:1992]. How listeners experience digital audio depends on sound fidelity, i.e. the accuracy of a digital signal representing the reproduction of sound. High-fidelity or hi-fi digital audio ideally should have low levels of noise and distortion in its signal to provide a high perceptual quality.
Our goal is to assess widely used audio compression codecs such as [FLACWebsite] or AAC and MP3 [Brandenburg:MP3AACE:1999] with respect to compression efficiency as well as how the compression techniques in these audio codecs affect the perceptual quality of the signal reconstruction. In addition, we use visual analytics derived techniques to present the perceptual impact of compression techniques via spectrograms, sound field diagrams, and objective perceptual scoring using Perceptual Evaluation of Audio Quality (PEAQ) [Thiede:PEAQITUSOMPAQ:2000].
We argue that in choosing an audio compression codec, users should not only look at the compression efficiency but also consider the sonic perceptual quality of an audio compression codec.
2 Related Work
Audio codecs use general compression techniques to reduce the size of digital audio file. Traditional audio compression techniques include Huffman coding [Bosi:IDACS:2003], Linear Predictive Coding (LPC) [Liebchen:IMALC:2004], and psychoacoustic-based perceptual encoding [Burg:DSMCAS:2016].
Audio codecs are categorized into lossy or lossless depending on the compression scheme. Lossy audio codecs, such as MP3 and AAC [Brandenburg:MP3AACE:1999] or Vorbis [Moffitt:OVOFA:2001], compress digital audio by removing unimportant data, i.e. parts of the digital signal that can be difficult to perceive by a listener because it is not within the human hearing threshold. Lossless audio codecs such as [FLACWebsite] compress digital audio by only removing redundancies and compressing repeated patterns in the audio signal while maintaining all of the original signal’s data.
Assessing digital audio codecs requires comparing compression efficiency as well as evaluating the resulting perceptual quality. Compression metrics include compression ratio, encoding speed, and decoding speed. Evaluating perceptual quality requires a more complex assessment. Classically, perceptual quality assessments require trained listeners to perform listening tests [Olive:HHTL:2011]. However, listening tests are a rather subjective assessment with inherent listener bias as well as reproducibility problems. We present perceptual-quality assessment using visualizations as well as objective scoring. Tools such as frequency spectra [AudacityWebsite], spectrograms [SpekWebsite], and sound-field diagrams [IzotopeWebsite] show the impact of compression artifacts to the perceptual quality of encoded digital audio. In addition, Perceptual Evaluation of Audio Quality (PEAQ) is used to quantify the perceptual quality by generating a single Objective Difference Grade (ODG) score of encoded digital audio [Thiede:PEAQITUSOMPAQ:2000].
In recent years, experiments with AI-driven audio compression techniques have been conducted. These techniques are based on well-known neural network architectures such as Residual Vector Quantized Generative Adversarial Networks (RVQGAN) [Kumar:HFACIRVQGAN:2023]. We also compare RVQGAN audio compression technique with traditional audio compression techniques in our evaluation.
3 Experimental Setup
3.1 Hardware and Software
All experiments were conducted on a computer workstation featuring an AMD Ryzen 5 5600X CPU, 32GB RAM at 3200 MHz, and a Samsung 970 EVO NVMe SSD storage device. The reference listening setup included a [DangerousMusicWebsite] digital-to-analog converter (DAC), a [SangakuDIYWebsite], and [SennheiserHD600Website] open-back headphones.
Encoding and decoding of digital audio used [FMediaWebsite], a flexible audio processing tool supporting most existing lossy and lossless codecs. Additionally, fmedia also allowed recording of file sizes as well as elapsed time for audio encoding and audio decoding processes. These measurements were used to calculate compression ratio, audio encoding speed, and audio decoding speed.
3.2 Audio Codecs and Settings
Audio-compression ratio in percentage is calculated as shown in equation 1.
$\text{Compression ratio}=\frac{\text{Uncompressed file size}}{\text{Compressed file size}}\cdot 100$ (1)
The lower the compression ratio, the more storage space a digital audio file will occupy. Encoding speed and decoding speed are expressed in samples per microsecond and calculated as shown in equation 2.
$\text{Speed}=\frac{2\cdot\text{Sample size}}{\text{Recorded time}\cdot 1,000,000}$ (2)
Sample size in equation 2 is multiplied by two because audio tracks (usually) contain stereo signals in a left and a right channel. In addition, calculated speeds are divided by 1,000,000 in equation 2 to convert from samples per second to samples per microseconds.
The most commonly used audio codecs were chosen for our experiments: FLAC, MP3, AAC, and Vorbis. These audio codecs have different settings, allowing users to choose different compression levels. To keep our experiments close to real life scenarios, the most common settings were chosen for each audio codec. Table 1 provides an overview of the settings used for the various audio encoding formats.
Codec Option
Name Range Setting
Lossless FLAC Compression 0 – 8 6
MP3 CBR 64 – 320 kbps 320 kbps
MP3 CBR 64 – 320 kbps 128 kbps
Lossy AAC CBR 8 – 800 kbps 256 kbps
AAC VBR 1 – 5 5
Vorbis VBR 1 – 10 7
Table 1: Audio-codec settings. The Range column shows an option’s range while the Settings column shows the used setting. CBR stands for constant bit rate, VBR stands for variable bit rate.
3.3 Visualization Tools
To visualize perceptual quality, a combination of spectral analysis tools were employed. These include built-in spectrum generator from [AudacityWebsite] (cf. figure 1a), spectrogram analysis by [SpekWebsite] (cf. figure 1b), and sound-field visualizations provided by [IzotopeWebsite] (cf. figure 1c).
(a) Spectrum
(b) Spectrogram
(c) Sound field
Figure 1: Visualization of perceptual audio quality. LABEL:sub@fig:setup:visual:spectrum+sample Audacity generated spectrum shows peak loudness within the highest frequency (48 kHz); x-axis shows time, y-axis shows frequency, colored spectrum indicates the loudness in dB. LABEL:sub@fig:setup:visual:spectrogram Spek generated spectrogram shows loudness range (0 dB to -120 dB) of audio signal’s frequency range; x-axis shows time, y-axis shows frequency range, colored spectrum indicates loudness in dB. LABEL:sub@fig:setup:visual:sound+field Sound field visualization generated by Audacity using the Insight 2 plug-in shows width, depth, and height of the sound stage as well as the location of individual sound sources within the sound stage. The half-circle line indicates front of the sound stage while bright dots indicate stereo images. All visualizations based on Dreams by Fleetwood Mac (2001 remaster vinyl quality, 96 kHz sample rate).
Each type of visualization represents a different aspect of aural perceptual quality. Spectrum analysis shows how audio compression techniques affect the loudness of a digital audio file through its frequency response. Spectrograms display time, frequency, and amplitude with a color range for easy visual differentiation. Sound fields reveals how audio compression techniques affect the stereo image and sound stage, i.e. how audio signals are distributed between the left and the right channel of a stereo system and the width, depth and height of an audio signal, respectively.
3.4 Perceptual Evaluation of Audio Quality
To quantify the impact of audio-compression techniques on digital audio signals, an objective scoring system is required. Perceptual Evaluation of Audio Quality (PEAQ) [Thiede:PEAQITUSOMPAQ:2000] is an ITU-standardized algorithm to assess the perceptual quality of digital audio. PEAQ generates an ODG score (cf. table 2), expressed in terms of a quality gap between a tested audio signal and its original audio signal.
Impairment ODG
Imperceptible 0
Perceptible, not annoying -1
Slightly annoying -2
Annoying -3
Very annoying -4
Table 2: Objective Difference Grade scoring range, after [ITU:BS.1387-2:2023]. A generated score from the objective measurement corresponds to a Subjective Difference Grade (left column). Resolution of ODG scores is limited to one decimal.
Advanced PEAQ (APEAQ) [Kabal:EIITUPEAQ:2002] is a more appropriate version of PEAQ for newer audio codecs [Delgado:CWSUP:2020], while basic PEAQ’s total Noise-to-Mask (totalNMR) model output variable (MOV) is well-correlated with subjective listening tests [Dick:ODAQ:2025]. In our study, we used both APEAQ’s ODG and basic PEAQ’s totalNMR MOV as objective perceptual quality evaluation metrics. totalNMR is measured in negative decibels and, ideally, should be as far as possible from 0 dB while at the same time totalNMR is trending close to 0 dB.
To calculate APEAQ’s ODG and PEAQ’s totalNMR we used GstPEAQ [Holters:GSTPEAQ:2015] and [peaqbGitHub], respectively. It is important to note that these open-source projects do not fully conform to [ITU:BS.1387-2:2023]. However, our pre-study experiments showed that both GstPEAQ as well as peaqb perform within reasonably close limits to each other and should not do worse than subjective listening test [Holters:GSTPEAQ:2015].
3.5 Test Samples
We conducted our experiments on 50 digital audio files, using high-fidelity conversions of CDs, DVDs, vinyls, or cassettes from a wide range of music genres and epochs. For each sample, we collected the file size, the sampling rate, and the original bit rate used for conversion. Each sample was then encoded using our chosen audio codecs at their respective quality settings (cf. table 1). To showcase the impact of audio-compression techniques on the perceptual quality of digital audio, we chose audio files for visualizations that present high levels of perceptual quality as well as a high degree of accuracy in stereo imaging and sound-stage appearance.
4 Evaluation
4.1 Audio-Codec Performance
Table 3 provides excerpts from the dataset for selected entries along with the most common quality setting, compression ratio as a percentage, and encoding speed as well as decoding speed in samples per microsecond.
Track Codec Setting Compr. Ratio Enc. Speed Dec. Speed
Dreams Fleetwood Mac FLAC level 6 68.17 37.91 117.93
MP3 320 kbps 89.59 5.92 158.79
MP3 128 kbps 95.84 7.20 210.69
AAC 256 kbps 91.57 8.86 37.86
AAC level 5 89.50 6.83 32.99
Vorbis level 7 92.50 9.47 58.46
La mer II. Jeux de Vagues Claude Debussy FLAC level 6 66.67 38.87 118.76
MP3 320 kbps 77.32 3.14 75.12
MP3 128 kbps 90.93 3.67 99.48
AAC 256 kbps 81.75 5.43 28.81
AAC level 5 84.16 5.01 29.26
Vorbis level 7 87.77 9.09 46.43
Foreigner Pallbearer FLAC level 6 42.48 36.41 104.45
MP3 320 kbps 77.33 3.64 75.80
MP3 128 kbps 90.93 3.91 96.80
AAC 256 kbps 81.76 5.58 28.43
AAC level 5 84.41 5.41 29.33
Vorbis level 7 86.70 8.36 44.04
September Earth, Wind & Fire FLAC level 6 29.98 26.33 82.76
MP3 320 kbps 93.06 6.02 148.92
MP3 128 kbps 97.22 8.10 207.89
AAC 256 kbps 94.38 9.06 38.02
AAC level 5 92.51 6.64 34.07
Vorbis level 7 94.48 9.09 56.52
Table 3: Excerpt from the codec-performance dataset showing audio codec compression as percentage ratio and encoding as well as decoding speeds in samples per microsecond using the audio codec configurations established in table 1.
The full dataset, [FMediaACCDGithubRepo], currently contains 53 entries with each entry also specifying additional metadata (e.g., genre) as well as its file size and the number of samples. This is followed by compression bit rate, the compressed file size, the compression, and the amount of time as well as the speed for encoding and decoding for each of the tested audio codecs.
Table 4 presents the average compression ratio, encoding speed and decoding speed of the dataset.
Codec Setting Comp. Ratio Enc. Speed Dec. Speed
Lossless FLAC level 6 38.74 31.92 92.47
Lossy MP3 320 kbps 83.99 4.90 112.30
MP3 128 kbps 92.44 5.70 148.93
AAC 256 kbps 86.73 7.12 33.07
AAC level 5 87.42 6.03 31.97
Vorbis level 7 88.68 8.69 46.39
Table 4: Average audio-codec compression as percent ratio and encoding as well as decoding speeds in samples per microsecond. Best performance in each category has been marked.
FLAC, the only lossless compression scheme tested, achieves an average encoding speed of 32 samples per microsecond, i.e. 32 million samples per second. MP3 at 128 kbps, a lossy compression scheme, exhibits both the best average compression ratio at 92% as well as the best decoding speed at 149 samples per microsecond, i.e. 149 million samples per second.
4.2 Perceptual Quality
Figure 2 visualizes the spectra of all tested audio codecs versus the uncompressed digital audio file of the song Iron Man by Black Sabbath.
Figure 2: Spectra for all tested audio codecs (exported from Audacity). X-axis denotes the frequency range up to 24 kHz. The y-axis shows the loudness range from 0 dB to -150 dB. The color-coded graphs refer to uncompressed (black), FLAC level 6 (red), MP3 CBR 320 kbps (yellow), MP3 CBR 128 kbps (green), AAC CBR 256 kbps (orange), AAC VBR level 5 (teal), and Vorbis VBR level 7 (pink). The graphs represent encoded audio signal outputs for the respective audio codec. Note the spectrum for the uncompressed audio signal is completely aligned with the spectrum of FLAC level 6, showing that FLAC encoded audio exhibits the exact same loudness as the uncompressed audio signal across the entire frequency range.
The graph of FLAC level 6 aligns perfectly with the spectrum of uncompressed audio, showing that the loudness of FLAC encoded audio exactly matches the loudness of the uncompressed audio across the entire spectrum. On the other hand, spectra for all audio encoded with lossy compression techniques deviate from the spectrum of the original audio signal. The same pattern of reduction in loudness can be seen for the interval 15 kHz to above 20 kHz. MP3 320 kbps and MP3 128 kbps elevate the loudness of the entire frequency range before cutoff at 18 kHz and 15 kHz, respectively. AAC CBR 256 kbps loudness reduction happens much earlier in the frequency range, around 7 kHz. To our surprise, AAC VBR level 5 loudness closely matches with the spectra for FLAC and the uncompressed audio very well with only a slight loudness reduction at the high-frequency range. Vorbis level 7 also exhibits loudness reduction at high frequencies. However, unlike the other tested lossy compression techniques, it manifests a smooth decline in loudness.
[SpekWebsite] was used to further investigate and visualize the loudness change in encoded audio compared to uncompressed audio. Figure 3 shows that all audio encoded with lossy compression schemes exhibit very quiet loudness in the range -100 dB to -105 dB, marked dark blue in figure 3a, which is further reduced to near silence at -120 dB, marked black in figures 3c to 3g, for high frequencies above 15 kHz.
Figure 3: Audible loudness comparison for the track Iron Man by Black Sabbath using different encoding schemes. LABEL:sub@fig:evaluation:iron+man:a shows the spectrogram of the uncompressed signal while LABEL:sub@fig:evaluation:iron+man:b – LABEL:sub@fig:evaluation:iron+man:g show spectrograms for FLAC, MP3, AAC, and Vorbis, respectively. Note that MP3 requires a version of the track scaled down from 96 kHz to 48 kHz; the diagram for MP3 has been scaled in turn by a factor of two. All x-axes show duration of the tested audio signal while y-axes show the frequency range of the tested audio signal. Spectrograms indicate loudness in dB.
MP3 320 kbps retains some portion of very quiet loudness until 20 kHz, while AAC VBR level 5 retains some very quiet loudness. Vorbis level 7 loudness reduction is not as flat as other lossy codecs.
Figure 4 shows example sound-field visualizations for stereo images and sound stages capturing the 3:00 – 3:30 min interval of Iron Man by Black Sabbath, when Tony Iommi performs the guitar solo.
Figure 4: Sound-field visualizations for the track Iron Man by Black Sabbath compressed using different audio encoding schemes. LABEL:sub@fig:evaluation:iron+man+sf:a shows the sound-field visualization from 3:00 – 3:30 min for the uncompressed signal and LABEL:sub@fig:evaluation:iron+man+sf:b – LABEL:sub@fig:evaluation:iron+man+sf:g show the sound-field visualization for the same time interval encoded with FLAC, MP3s, AACs, and Vorbis, respectively. The patterns of dots should lean to the right, capturing the stereo image of the electric guitar’s sound in the recording. Compared to the uncompressed audio signal’s stereo image, the FLAC encoded audio signal was able to accurately capture the stereo image while audio signals from other lossy encoders show some inaccuracies.
Listening to the song, the guitar solo sounds slightly to the right, so the stereo image of this session should lean to the right of the center as well. All the sound-field visualizations in figures 4b to 4g have their clusters lean to the correct direction. However, the clusters in figures 4c to 4g, representing audio encoded with lossy compression techniques, significantly deviate from the uncompressed sound field in figure 4a. The sound-field clusters are too sparse for figures 4c to 4d using MP3 encoding, while the clustering for AAC CBR 256 kbps in figure 4e is too condensed in the center. However, both AAC VBR level 5 (cf. figure 4f) and Vorbis level 7 (cf. figure 4g) sound-field clusters are very similar to the uncompressed audio sound field in figure 4a.
Finally, advanced PEAQ ODG scores and totalNMR were calculated based on GstPEAQ [Holters:GSTPEAQ:2015] and [peaqbGitHub], respectively, to quantify the impact of compressing digital audio signals on perceptual quality. Table 5 provides advanced PEAQ ODG scores and totalNMR values for all tested samples in table 3. Advanced PEAQ ODG score range is 0 to -4 table 2 and the lower totalNMR the better.
Track Codec Setting Adv. ODG totalNMR
Dreams Fleetwood Mac FLAC level 6 0.2 -131.08
MP3 320 kbps -0.1 -0.77
MP3 128 kbps -0.1 -1.07
AAC 256 kbps -0.1 0.15
AAC level 5 -0.1 0.15
Vorbis level 7 -0.1 -16.75
La mer II. Jeux de Vagues Claude Debussy FLAC level 6 0.2 -121.33
MP3 320 kbps -2.0 -2.04
MP3 128 kbps -2.4 -2.4
AAC 256 kbps -1.4 -1.03
AAC level 5 -1.4 -1.03
Vorbis level 7 -0.1 -17.56
Foreigner Pallbearer FLAC level 6 0.2 -126.49
MP3 320 kbps -0.6 -3.04
MP3 128 kbps -0.2 -3.33
AAC 256 kbps -0.2 -2.35
AAC level 5 -0.2 -2.34
Vorbis level 7 -0.2 -17.28
September Earth, Wind & Fire FLAC level 6 0.2 -135.91
MP3 320 kbps -0.2 -1.14
MP3 128 kbps -3.5 -1.43
AAC 256 kbps -0.2 -0.46
AAC level 5 -0.2 -0.46
Vorbis level 7 -0.1 -16.74
Table 5: Advanced PEAQ ODG score and totalNMR for all tested samples in table 3.
Table 6 presents the average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs at their most common quality settings.
Codec Setting Adv. ODG totalNMR
FLAC level 6 0.2 -130.11
MP3 CBR 320 kbps -0.2 -1.34
MP3 CBR 128 kbps -0.4 -1.72
AAC CBR 256 kbps -0.4 -0.08
AAC VBR level 5 -0.4 -0.08
Vorbis VBR level 7 -0.1 -16.74
Table 6: Average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs. While all audio codecs score near 0 on advanced PEAQ ODG, FLAC’s average totalNMR is significantly lower than the average totalNMR of all lossy codecs, with Vorbis comes in second.
All audio codecs score near 0 on average advanced PEAQ ODG score. However, FLAC’s average totalNMR is the lowest. Vorbis’s average totalNMR is second lowest but there is a significant gap between FLAC and Vorbis. The remaining average totalNMR are -1 to 0.
4.3 Machine-Learning Based Codec
In recent years, a trend emerged to apply machine-learning techniques into remastering and compressing audio. One proposed audio compression model employs a residual vector quantized generative adversarial network (RVQGAN) [Kumar:HFACIRVQGAN:2023]. RVQGAN works by encoding the audio signal into a lower-dimensional presentation followed by converting this presentation into discrete values. We employed RVQGAN’s prototype implementation from its [DACRVQGANGithubRepo] and encoded the track Iron Man by Black Sabbath.
Table 7 presents RVQGAN’s average compression ratio, average encoding speed and average decoding speed together with traditional audio codecs for comparison.
Codec Setting Comp. Ratio Enc. Speed Dec. Speed
Lossless FLAC level 6 38.74 31.92 92.47
Lossy MP3 320 kbps 83.99 4.90 112.30
MP3 128 kbps 92.44 5.70 148.93
AAC 256 kbps 86.73 7.12 33.07
AAC level 5 87.42 6.03 31.97
Vorbis level 7 88.68 8.69 46.39
AI RVQGAN 98.15 1.45 1.03
Table 7: Average audio-codec compression as percent ratio and encoding as well as decoding speeds in samples per microsecond. Values for lossless and lossy encoding are the same as in table 4 plus the results for RVQGAN. Best performance in each category has been marked.
In its current state RVQGAN does improve on the compression ratio by reducing storage size to almost 1% of the original size. However, encoding and decoding speed are nearly one to two magnitudes slower than any of the traditional audio codecs.
Figures 5 to 7 illustrate the spectrum, spectrogram, and sound-field visualization of an RVQGAN encoded audio signal compared to the original uncompressed audio signal.
Figure 5: RVQGAN’s spectrum (red) together with the uncompressed signals spectrum (blue) (exported from Audacity). X-axis shows the frequency range of the sample track in Hz, y-axis shows loudness range from 0 dB to -125 dB. Compared to the uncompressed audio signal, the RVQGAN encoded audio signal shows a slight increase in loudness before dropping off at around 10 kHz, gradually reducing in volume towards higher frequencies.
Figure 6: Audible loudness comparison for the track Iron Man by Black Sabbath compressed using LABEL:sub@fig:evaluation:iron+man:rvqgan RVQGAN. Compared to the uncompressed audio signal in LABEL:sub@fig:evaluation:iron+man:og, the RVQGAN encoded audio signal exhibits a cut-off in loudness at around 21 kHz. X-axis shows duration of the tested audio signal, y-axis shows the frequency range of the tested audio signal. Color spectrum indicates loudness in dB.
Figure 7: Sound-field visualizations at 3:00 – 3:30 min for the track Iron Man by Black Sabbath. LABEL:sub@fig:evaluation:iron+man:og+sf uncompressed recording and LABEL:sub@fig:evaluation:iron+man:rvqgan+sf compressed with RVQGAN. It is apparent that the pattern in the RVQGAN encoded audio signal output is more smeared and lacks the accuracy of the guitar’s position when compared to the uncompressed audio’s sound field.
Despite achieving an impressive compression ratio of around 98% of the uncompressed file size, perceptual quality of RVQGAN encoded audio is questionable. Currently, RVQGAN exhibits worse perceptual quality than all traditional audio codecs. RVQGAN’s average ODG score and average totalNMR in table 8 supports this evaluation.
Codec Setting Adv. ODG totalNMR
FLAC level 6 0.2 -130.11
MP3 CBR 320 kbps -0.2 -1.34
MP3 CBR 128 kbps -0.4 -1.72
AAC CBR 256 kbps -0.4 -0.08
AAC VBR level 5 -0.4 -0.08
Vorbis VBR level 7 -0.1 -16.74
RVQGAN -3.7 -0.69
Table 8: Average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs. This is table 6 with the additional advanced PEAQ ODG score and totalNMR for RVQGAN.
5 Discussion
5.1 Audio-Codec Performance
In our audio-codec performance tests (cf. table 4), FLAC exhibited the lowest compression ratio of all tested audio codecs, meaning that lossless audio files will take up more storage space than files encoded using a lossy compression technique. On the other hand, FLAC also achieved the highest encoding speed. These two average values for FLAC are expected because lossless compression techniques preserve most data in an audio file while only removing redundancies and repeated patterns.
All lossy audio codecs have similar compression ratio, with MP3 (128 kbps) achieving the highest compression ratio. Both MP3 quality settings exhibit lower encoding speeds than other lossy audio codecs. On the other hand, they show significantly higher decoding speeds than other lossy audio codecs with MP3 at 128 kbps being the fastest to decode. It can be argued that MP3 is a strong contender for the best lossy audio codec due to its high decoding speeds, making it the ideal choice for audio playback on low-power devices.
Among lossy audio encoders, Vorbis exhibits above average compression ratios, the fastest encoding speeds, and decent decoding speeds. In our audio-codec performance evaluation, Vorbis is the runner up for lossy audio-codecs, providing decent compression ratios and fastest lossy encoding speeds.
5.2 Perceptual Quality
Our perceptual quality evaluation shows that lossy compression algorithms sacrifice high-frequency aural details to reduce storage size. Audio encoding using psycho-acoustic principles determines these region are mostly inaudible to the average listener and the audio data for these regions can simply be removed. The differing graphs in figure 2 for all lossy audio codecs lies in how each codec handles loudness reduction of the high-frequency region.
Loudness reduction or removal of high frequencies reduces the harmonic richness, the details and clarity, and the spatial depth [Truax:THFAE:2020] that allow digital music to exhibit high fidelity audio. [Vickers:TLWBSR:2010], mainly reporting on the loudness war in music production, also shows that the average human listener can perceive aural quality very differently depending on the loudness alteration in a music recording. Additionally, loudness reduction and high-frequency removals affect the stereo image of the encoded audio reducing their aural accuracy.
The APEAQ ODG scores and totalNMR values we collected from all tested audio compression codecs (cf. table 6) support our findings. On average, all traditional audio codecs - FLAC, MP3, AAC, and Vorbis - score near 0 on APEAQ ODG. This means that APEAQ evaluates all of these traditional audio codecs impairment level as imperceptible (cf. table 2). However, the noise level of lossy codecs is higher than noise level of lossless codecs. In particular, noise levels for MP3 and AAC are right at the threshold to their mask levels, meaning their noise levels are potentially perceptible. Remarkably, Vorbis is the only lossy audio-compression codec where the noise level is masked reasonably well, making it the winner of this category.
While most of our assessment was conducted via a system of codec performance measurements, visualizations, and PEAQ tests, we have to address the importance of listening tests. PEAQ were built under [ITU:BS.1387-2:2023] to predict how an average human perceive audio quality [ITU:BS.1387-2:2023]. Trained listeners, under [ITU:BS.1534-3:2015] on conducting MUSHRA tests, need to meet certain requirements to be qualified as assessors [ITU:BS.1534-3:2015]. Thus, they can provide important and practical feedback about how humans perceive aural quality, which validate the result of the objective evaluation models, hence validate our study.
5.3 AI-based Audio Codecs
Regarding AI-based compression technique, it is known that there were several neural-network models that were proposed. These models include an end-to-end approach [Rim:DNNETE:2021], lightweight transformer approach [Défossez:HFNAC:2022], and a novel audio codec called APCodec [Ai:APCodec:2024]. We chose the RVQGAN model by [Kumar:HFACIRVQGAN:2023] because the authors claimed their model outperforms competing traditional audio compression algorithms.
Our tests indicate RVQGAN indeed exhibits the highest compression ratio compared with traditional audio codecs (cf. table 7). Storage size of an RVQGAN encoded audio signal is only around 1% to 2% of the size of the uncompressed signal, which is an impressive result. However, RVQGAN’s encoding and decoding speeds are very low compared to traditional audio codecs, performing one to two magnitudes slower. Additionally, in its current version, RVQGAN’s output is a unique format descript-audio-codec (.dac) that is not playback-supported by any audio player application at the time of our evaluation tests.
While the RVQGAN codec achieved the highest average compression ratio in our audio-codec performance tests, Our testing shows that RVQGAN’s perceptual quality is worse than traditional audio codecs. RVQGAN’s average advanced ODG score is -3.7 which is close to very annoying (cf. table 2), and RVQGAN’s average totalNMR indicates that its noise level is quite perceptible (cv. table 8). The AI-based approach to developing audio codecs shows promise in advancing efficiency but will require further refinements to provide acceptable hi-fi listening experiences to end users.
6 Conclusions
We evaluated several of the most common audio-compression codecs providing insights in how audio compression techniques affect sonic perceptual quality of digital music. Our tests indicate that lossy audio-compression codecs do provide better compression performance than lossless audio compression codecs. This stems mainly from the fact that lossy audio-compression algorithms sacrifice aural perceptual quality for higher compression speeds and lower final storage size. However, the Vorbis audio codec emerged as the exception to this rule. It exhibits acceptable compression efficiency when compared to its competitors while at the same time achieving perceptual quality nearly imperceptibly different compared to respective uncompressed digital audio signal.
We also evaluated RVQGAN, an AI-based audio compression scheme. While RVQGAN’s codec performance in speed and size was impressive, its perceptual quality is severely lacking. We are sure to see more advancement in this field, especially improving the perceptual quality of future AI-driven audio compression codec developments.
Our evaluation strengthens our argument that, when choosing a digital audio compression codec, focus should not only be on the audio compression performance. Users should also consider the perceptual quality of the encoded audio signal to ensure the best possible listening experience. We compliment Vorbis as an excellent lossy audio-compression codec also able to maintain high-fidelity perceptual quality.
6.1 Future Work
Expanding the database of audio codec performance measurements can determine the consistency in compression ratio, encoding speed, and decoding speed of all audio compression codec. We welcome readers to add performance measurements to our database https://github.com/samcate/fmedia-Audio-Codec-Comparison-Database.
With the recent trend of high-fidelity music streaming services like [Spotify], perceptual quality evaluation should be conducted on streaming audio. Streaming services utilize a wide range of audio-compression schemes, which can provide interesting insights.
Future research should also focus on enhancing perceptual quality in AI-driven compression methods. By integrating psycho-acoustic principles into the compression neural network, it may be possible to achieve more accurate perceptual quality, providing low-storage size high-fidelity digital music to users.

### v2

@@
 Evaluation of Audio Compression Codecs
-abstract=true, fontsize=10pt, footinclude=false, headinclude=false, headsepline=false, mpinclude=false, numbers=noenddot, paper=letter, parskip=half-, twocolumn=true, twoside=false, [subfloat]aboveskip=0pt, belowskip=0pt, font=scriptsize,sf, format=hang, labelsep=space, margin=1pt, singlelinecheck=0,
-Evaluation of Audio Compression Codecs
-Thien T. Duong    Jan P. Springer
 Abstract
 Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity.
-It is how humans respond to the accuracy, intelligibility, and fidelity of aural media.
-Today this fidelity is also heavily influenced by the use of audio compression codecs for storing aural media in digital form.
-We argue that, when choosing an audio compression codec, users should not only look at compression efficiency but also consider the sonic perceptual quality properties of available audio compression codecs.
-We evaluate several commonly used audio compression codecs in terms of compression performance as well as their sonic perceptual quality via codec performance measurements, visualizations, and PEAQ scores.
-We demonstrate how perceptual quality is affected by digital audio compression techniques, providing insights for users in the process of choosing a digital audio compression scheme.
+It defines how humans respond to the accuracy, intelligibility, and fidelity of aural media.
+Today this fidelity is also heavily influenced by the use of audio-compression codecs used for storing aural media in digital form.
+We argue that, when choosing an audio-compression codec, users should not only look at compression efficiency but also consider the sonic perceptual quality properties of available audio-compression codecs.
+We evaluate several audio-compression codecs, traditional as well as AI-based, in terms of compression performance and their sonic perceptual quality via codec-performance measurements, various visualization techniques, and a family of PEAQ scores.
+We demonstrate how perceptual quality is affected by digital audio-compression techniques, providing insights for a wide range of users in the process of choosing a digital audio-compression scheme.
 Keywords: Audio Codecs, Audio Compression, Information Visualization, Quantitative and Qualitative Audio Evaluation, PEAQ
 1 Introduction
-Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity [Ellis:PRA:1992].
+Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity [undefl].
 How listeners experience digital audio depends on sound fidelity, i.e.
 the accuracy of a digital signal representing the reproduction of sound.
-High-fidelity or hi-fi digital audio ideally should have low levels of noise and distortion in its signal to provide a high perceptual quality.
-Our goal is to assess widely used audio compression codecs such as [FLACWebsite] or AAC and MP3 [Brandenburg:MP3AACE:1999] with respect to compression efficiency as well as how the compression techniques in these audio codecs affect the perceptual quality of the signal reconstruction.
-In addition, we use visual analytics derived techniques to present the perceptual impact of compression techniques via spectrograms, sound field diagrams, and objective perceptual scoring using Perceptual Evaluation of Audio Quality (PEAQ) [Thiede:PEAQITUSOMPAQ:2000].
-We argue that in choosing an audio compression codec, users should not only look at the compression efficiency but also consider the sonic perceptual quality of an audio compression codec.
+High-fidelity or hi-fi digital audio ideally should have low levels of noise and distortion in its signal to provide a high perceptual reproduction quality.
+Our goal is to assess widely used audio-compression codecs such as [undefm] or AAC and MP3 [undefc] as well as AI-based audio codecs such as RVQGAN [undefw] with respect to compression efficiency as well as how the compression techniques in these audio codecs affect the perceptual quality of the signal reconstruction.
+We use visual analytics derived techniques to present the perceptual impact of compression techniques via spectrograms, sound-field diagrams, and objective perceptual scoring based on Perceptual Evaluation of Audio Quality (PEAQ) [undefaf] and its variant implementations [undefu].
+While subjective listening tests are considered the gold standard for audio-quality evaluation, our test method is based on objective quality scoring using PEAQ.
+We believe this will allow for more straightforward replication by non-domain experts as well as allowing additional results to be added into community maintained repositories.
+We argue that in choosing an audio-compression codec, users should not only look at the compression efficiency but also consider the sonic perceptual quality of an audio-compression codec in reproducing audio signals.
 2 Related Work
-Audio codecs use general compression techniques to reduce the size of digital audio file.
-Traditional audio compression techniques include Huffman coding [Bosi:IDACS:2003], Linear Predictive Coding (LPC) [Liebchen:IMALC:2004], and psychoacoustic-based perceptual encoding [Burg:DSMCAS:2016].
-Audio codecs are categorized into lossy or lossless depending on the compression scheme.
-Lossy audio codecs, such as MP3 and AAC [Brandenburg:MP3AACE:1999] or Vorbis [Moffitt:OVOFA:2001], compress digital audio by removing unimportant data, i.e.
+Audio codecs use general compression techniques to reduce the size of digital-audio content.
+Traditional audio compression techniques include Huffman coding [undefq, undefb], Linear Predictive Coding (LPC) [undefx], and psychoacoustic-based perceptual encoding [undefd].
+Audio codecs are categorized into lossy and lossless depending on the employed compression scheme.
+Lossy audio codecs, such as MP3 and AAC [undefc] or Vorbis [undefy], compress digital audio by removing unimportant data, i.e.
 parts of the digital signal that can be difficult to perceive by a listener because it is not within the human hearing threshold.
-Lossless audio codecs such as [FLACWebsite] compress digital audio by only removing redundancies and compressing repeated patterns in the audio signal while maintaining all of the original signal’s data.
+Lossless audio codecs such as [undefm] encode digital audio by only removing redundancies and compressing repeated patterns in the audio signal while maintaining all of the original signal’s data.
 Assessing digital audio codecs requires comparing compression efficiency as well as evaluating the resulting perceptual quality.
 Compression metrics include compression ratio, encoding speed, and decoding speed.
-Evaluating perceptual quality requires a more complex assessment.
-Classically, perceptual quality assessments require trained listeners to perform listening tests [Olive:HHTL:2011].
-However, listening tests are a rather subjective assessment with inherent listener bias as well as reproducibility problems.
-We present perceptual-quality assessment using visualizations as well as objective scoring.
-Tools such as frequency spectra [AudacityWebsite], spectrograms [SpekWebsite], and sound-field diagrams [IzotopeWebsite] show the impact of compression artifacts to the perceptual quality of encoded digital audio.
-In addition, Perceptual Evaluation of Audio Quality (PEAQ) is used to quantify the perceptual quality by generating a single Objective Difference Grade (ODG) score of encoded digital audio [Thiede:PEAQITUSOMPAQ:2000].
-In recent years, experiments with AI-driven audio compression techniques have been conducted.
-These techniques are based on well-known neural network architectures such as Residual Vector Quantized Generative Adversarial Networks (RVQGAN) [Kumar:HFACIRVQGAN:2023].
-We also compare RVQGAN audio compression technique with traditional audio compression techniques in our evaluation.
+Evaluating perceptual quality requires more complex assessments.
+Classically, perceptual-quality assessments require trained listeners to perform listening tests [undefz], which is considered the gold standard in perceptual-quality evaluation.
+A subjective listening test such as MUSHRA [undefs] requires trained listeners to be accurate.
+However, it has been proven that biases exist in MUSHRA [undefaj].
+Additionally, recent media coverage of blind listening tests has brought attention to instances where listeners were unable to discriminate between sound transmitted through significantly different types of conductive materials [undeff].
+This emphasizes our view that subjective listening tests do not necessarily match human intuition regarding audio quality.
+We present our perceptual-quality assessment using various visualization techniques as well as objective scoring to evaluate model-predicted perceptual impact.
+Tools such as frequency spectra [undefa], spectrograms [undefad], and sound-field diagrams [undeft] allow exposing the impact of compression artifacts on the perceptual quality of compressed digital audio.
+Additionally, Perceptual Evaluation of Audio Quality (PEAQ) [undefaf] is used to quantify the perceptual quality of audio signals by generating a single Objective Difference Grade (ODG) score of encoded digital audio.
+However, PEAQ is not perfect and there are many attempts to improve on the objective evaluation model.
+Advanced PEAQ [undefu] has been suggested for modern audio codecs, which tend to exhibit better perceptual encoding algorithms.
+[undefk] show that by using the PEAQ’s model output variable (MOV) totalNMR by itself is already a good evaluation metric in combination with advanced PEAQ’s ODG score.
+Additionally, the 2f model [undefv] in particular is a good example for objectively approximating subjective MUSHRA scores [undefj].
+In recent years, experiments with audio compression techniques based on machine-learning and deep-learning techniques have been conducted.
+There have been several AI-based compression technique proposed, which include end-to-end approaches [undefaa], lightweight-transformer approaches [undefi], and novel audio codecs such as APCodec [undef].
+We chose [undefw]’s Residual Vector Quantized Generative Adversarial Networks (RVQGAN) model because the authors’ claim that their model outperforms competing traditional audio compression algorithms.
+We do not intend to replace subjective listener based evaluation with our work.
+Rather we explore how computational-perceptual models such as PEAQ and its variants in combination with a variety of visualization techniques can facilitate reproducible benchmarking and exploratory evaluation in line with psychoacoustic principles for scalable audio-codec analysis.
 3 Experimental Setup
@@
 All experiments were conducted on a computer workstation featuring an AMD Ryzen 5 5600X CPU, 32GB RAM at 3200 MHz, and a Samsung 970 EVO NVMe SSD storage device.
-The reference listening setup included a [DangerousMusicWebsite] digital-to-analog converter (DAC), a [SangakuDIYWebsite], and [SennheiserHD600Website] open-back headphones.
-Encoding and decoding of digital audio used [FMediaWebsite], a flexible audio processing tool supporting most existing lossy and lossless codecs.
-Additionally, fmedia also allowed recording of file sizes as well as elapsed time for audio encoding and audio decoding processes.
+The reference listening setup included a [undefh] digital-to-analog converter (DAC), a [undefab], and [undefac] open-back headphones.
+Encoding and decoding of digital audio used [undefo], a flexible audio processing tool supporting most existing lossy and lossless codecs.
+Additionally, fmedia also allows recording of resulting file sizes as well as elapsed time for audio encoding and audio decoding processes.
 These measurements were used to calculate compression ratio, audio encoding speed, and audio decoding speed.
@@
 $\text{Compression ratio}=\frac{\text{Uncompressed file size}}{\text{Compressed file size}}\cdot 100$ (1)
-The lower the compression ratio, the more storage space a digital audio file will occupy.
+The lower the compression ratio the more storage space a digital-audio file will occupy.
 Encoding speed and decoding speed are expressed in samples per microsecond and calculated as shown in equation 2.
@@
 In addition, calculated speeds are divided by 1,000,000 in equation 2 to convert from samples per second to samples per microseconds.
-The most commonly used audio codecs were chosen for our experiments: FLAC, MP3, AAC, and Vorbis.
-These audio codecs have different settings, allowing users to choose different compression levels.
-To keep our experiments close to real life scenarios, the most common settings were chosen for each audio codec.
-Table 1 provides an overview of the settings used for the various audio encoding formats.
 Codec Option
 Name Range Setting
-Lossless FLAC Compression 0 – 8 6
+Lossless FLAC Compression 0 – 8 level 6
 MP3 CBR 64 – 320 kbps 320 kbps
@@
 Lossy AAC CBR 8 – 800 kbps 256 kbps
-AAC VBR 1 – 5 5
-Vorbis VBR 1 – 10 7
-Table 1: Audio-codec settings. The Range column shows an option’s range while the Settings column shows the used setting. CBR stands for constant bit rate, VBR stands for variable bit rate.
+AAC VBR 1 – 5 level 5
+Vorbis VBR 1 – 10 level 7
+Table 1: Default audio-codec settings for the audio codecs used in our tests.
+The Range column shows an option’s range while the Setting column displays the value used in testing.
+CBR stands for constant bit rate, VBR stands for variable bit rate.
+Instead of finding and using optimal parameters for each audio codec, presets were selected to represent frequently used default values.
+Our study prioritizes ecological validity and reproducibility over audio-codec tweaking when evaluating real-world audio-codec behavior under general use setups.
+Therefore, rather than describing theoretical audio-codec restrictions, reported results describe real-world deployment conditions.
+Table 1 provides an overview of the settings used for the various audio encoding formats.
 3.3 Visualization Tools
-To visualize perceptual quality, a combination of spectral analysis tools were employed.
-These include built-in spectrum generator from [AudacityWebsite] (cf.
-figure 1a), spectrogram analysis by [SpekWebsite] (cf.
-figure 1b), and sound-field visualizations provided by [IzotopeWebsite] (cf.
-figure 1c).
 (a) Spectrum
@@
 (c) Sound field
-Figure 1: Visualization of perceptual audio quality.
-LABEL:sub@fig:setup:visual:spectrum+sample Audacity generated spectrum shows peak loudness within the highest frequency (48 kHz); x-axis shows time, y-axis shows frequency, colored spectrum indicates the loudness in dB.
-LABEL:sub@fig:setup:visual:spectrogram Spek generated spectrogram shows loudness range (0 dB to -120 dB) of audio signal’s frequency range; x-axis shows time, y-axis shows frequency range, colored spectrum indicates loudness in dB.
-LABEL:sub@fig:setup:visual:sound+field Sound field visualization generated by Audacity using the Insight 2 plug-in shows width, depth, and height of the sound stage as well as the location of individual sound sources within the sound stage.
-The half-circle line indicates front of the sound stage while bright dots indicate stereo images.
-All visualizations based on Dreams by Fleetwood Mac (2001 remaster vinyl quality, 96 kHz sample rate).
+Figure 1: Perceptual-audio quality visualization.
+LABEL:sub@fig:setup:visual:spectrum+sample Audacity generated spectrum shows peak loudness within the highest frequency (48 kHz); x-axis shows time, y-axis shows frequency, colored spectrum indicates loudness in dB.
+LABEL:sub@fig:setup:visual:spectrogram Spek generated spectrogram shows loudness range (0 dB to -120 dB) of an audio signal’s frequency range; x-axis shows time, y-axis shows frequency range, colored spectrum indicates loudness in dB.
+LABEL:sub@fig:setup:visual:sound+field Sound-field visualization generated by Audacity using the Insight 2 plug-in shows width, depth, and height of the sound stage as well as the location of individual sound sources within that sound stage.
+The half-circle line indicates sound-stage front while bright dots indicate stereo images.
+Visualizations based on Dreams by Fleetwood Mac (2001 remaster vinyl quality, sampled at 96 kHz).
+To visualize perceptual quality a combination of spectral analysis tools were employed.
+These include the built-in spectrum generator of [undefa] (cf.
+figure 1a), spectrogram analysis by [undefad] (cf.
+figure 1b), and sound-field visualizations provided by [undeft] (cf.
+figure 1c).
 Each type of visualization represents a different aspect of aural perceptual quality.
-Spectrum analysis shows how audio compression techniques affect the loudness of a digital audio file through its frequency response.
+Spectrum analysis shows how audio-compression techniques affect the loudness of a digital-audio signal through its frequency response.
 Spectrograms display time, frequency, and amplitude with a color range for easy visual differentiation.
-Sound fields reveals how audio compression techniques affect the stereo image and sound stage, i.e.
-how audio signals are distributed between the left and the right channel of a stereo system and the width, depth and height of an audio signal, respectively.
+Sound fields reveal how audio-compression techniques affect the stereo image and the sound stage, i.e.
+how audio signals are distributed between the left and the right channel of a stereo system and the width, depth, and height of an audio signal, respectively.
 3.4 Perceptual Evaluation of Audio Quality
-To quantify the impact of audio-compression techniques on digital audio signals, an objective scoring system is required.
-Perceptual Evaluation of Audio Quality (PEAQ) [Thiede:PEAQITUSOMPAQ:2000] is an ITU-standardized algorithm to assess the perceptual quality of digital audio.
-PEAQ generates an ODG score (cf.
-table 2), expressed in terms of a quality gap between a tested audio signal and its original audio signal.
 Impairment ODG
@@
 Very annoying -4
-Table 2: Objective Difference Grade scoring range, after [ITU:BS.1387-2:2023].
-A generated score from the objective measurement corresponds to a Subjective Difference Grade (left column).
-Resolution of ODG scores is limited to one decimal.
-Advanced PEAQ (APEAQ) [Kabal:EIITUPEAQ:2002] is a more appropriate version of PEAQ for newer audio codecs [Delgado:CWSUP:2020], while basic PEAQ’s total Noise-to-Mask (totalNMR) model output variable (MOV) is well-correlated with subjective listening tests [Dick:ODAQ:2025].
-In our study, we used both APEAQ’s ODG and basic PEAQ’s totalNMR MOV as objective perceptual quality evaluation metrics.
-totalNMR is measured in negative decibels and, ideally, should be as far as possible from 0 dB while at the same time totalNMR is trending close to 0 dB.
-To calculate APEAQ’s ODG and PEAQ’s totalNMR we used GstPEAQ [Holters:GSTPEAQ:2015] and [peaqbGitHub], respectively.
-It is important to note that these open-source projects do not fully conform to [ITU:BS.1387-2:2023].
-However, our pre-study experiments showed that both GstPEAQ as well as peaqb perform within reasonably close limits to each other and should not do worse than subjective listening test [Holters:GSTPEAQ:2015].
+Table 2: Objective Difference Grade (ODG) scoring range, after [undefr].
+A generated score from the objective measurement corresponds to a Subjective Difference Grade (left column) qualitatively expressed as Impairment.
+ODG score resolution is limited to one decimal.
+To quantify the impact of audio-compression techniques on digital audio signals an objective scoring system is required.
+Perceptual Evaluation of Audio Quality (PEAQ) [undefaf] is an ITU-standardized algorithm to assess the perceptual quality of digital audio.
+PEAQ generates an Objective Difference Grade (ODG) score (cf.
+table 2), expressed in terms of a quality gap between a tested audio signal and its respective original audio signal.
+There are two versions of PEAQ: basic PEAQ (BPEAQ) and advanced PEAQ (APEAQ) [undefu].
+Both use the same Fast Fourier Transform (FFT) model to convert time-domain signals into frequency-domain signals for evaluation.
+APEAQ additionally includes a filter-bank ear model to accurately simulate the human auditory system’s cochlear response.
+[undefj] report BPEAQ to be less computational intensive than APEAQ, while APEAQ is the more appropriate version of PEAQ for newer audio codecs such as AAC and Vorbis.
+This is due to these audio codecs’ perceptual models being more complex.
+BPEAQ’s model output variable (MOV) total noise-to-mask ratio (totalNMR) is well-correlated with subjective listening tests [undefk].
+totalNMR is measured in negative decibels and ideally should be as far as possible from 0 dB while at the same time totalNMR is trending close to 0 dB.
+In our study, we calculate both basic and advanced PEAQ’s ODG scores as well as basic PEAQ’s totalNMR value to showcase the performance of each scoring metric.
+Additional models exist, which improve on PEAQ based on past evaluation.
+Of these the 2f model [undefv] performs significantly better than others [undefag].
+The 2f model uses the ADB and the AvgModDiff1 MOVs of the BPEAQ model to calculate a Mean MUSHRA Score Estimate ($\text{MMS}_{\text{est}}$) [undefv] as shown in equation 3.
+[…]
+The result of equation 3 is then clamped to the range [0…100] to calculate the actual 2f score as shown in equation 4.
+$\text{2f score}=\min(100,\max(0,\text{MMS}_{\text{est}}))$ (4)
+Past experiments have shown that the 2f model’s average score is the closest equivalent to a subjective listening test result [undefv].
+We use McGill University’s PEAQ Matlab model [undefu] to calculate a BPEAQ ODG score and extract the totalNMR MOV.
+In addition, ADB and AvgModDiff1 MOVs are extracted to calculate a 2f score.
+To calculate APEAQ’s ODG score, GstPEAQ [undefp] is utilized.
+We note that GstPEAQ is an open-source project that does not fully conform to [undefr].
+However, pre-study experiments have shown that GstPEAQ performs within reasonably close limit, i.e.
+0.03 difference in root mean square error, to [undefr] and is therefore expected to perform similar to a subjective listening test [undefp].
 3.5 Test Samples
-We conducted our experiments on 50 digital audio files, using high-fidelity conversions of CDs, DVDs, vinyls, or cassettes from a wide range of music genres and epochs.
+We conducted our experiments with 50+ digital-audio files, using high-fidelity conversions of CDs, DVDs, vinyls, and cassettes from a wide range of music genres and epochs.
 For each sample, we collected the file size, the sampling rate, and the original bit rate used for conversion.
@@
 table 1).
-To showcase the impact of audio-compression techniques on the perceptual quality of digital audio, we chose audio files for visualizations that present high levels of perceptual quality as well as a high degree of accuracy in stereo imaging and sound-stage appearance.
+To showcase the impact of audio-compression techniques on the perceptual quality of digital audio, in the following we chose a set of audio examples for visualizations that present high levels of perceptual quality as well as high degrees of accuracy in stereo imaging and sound-stage appearance.
 4 Evaluation
 4.1 Audio-Codec Performance
-Table 3 provides excerpts from the dataset for selected entries along with the most common quality setting, compression ratio as a percentage, and encoding speed as well as decoding speed in samples per microsecond.
-Track Codec Setting Compr. Ratio Enc. Speed Dec. Speed
+Track Codec Setting Compression Ratio Encoding Speed Decoding Speed
 Dreams Fleetwood Mac FLAC level 6 68.17 37.91 117.93
@@
 Vorbis level 7 94.48 9.09 56.52
-Table 3: Excerpt from the codec-performance dataset showing audio codec compression as percentage ratio and encoding as well as decoding speeds in samples per microsecond using the audio codec configurations established in table 1.
-The full dataset, [FMediaACCDGithubRepo], currently contains 53 entries with each entry also specifying additional metadata (e.g., genre) as well as its file size and the number of samples.
-This is followed by compression bit rate, the compressed file size, the compression, and the amount of time as well as the speed for encoding and decoding for each of the tested audio codecs.
-Table 4 presents the average compression ratio, encoding speed and decoding speed of the dataset.
-Codec Setting Comp. Ratio Enc. Speed Dec. Speed
+Table 3: Excerpt from our codec-performance dataset showing audio-codec compression as percentage ratio and encoding as well as decoding speeds in samples per microsecond using the audio-codec configurations established in table 1.
+Table 3 provides excerpts from our dataset for selected entries along with the chosen quality setting, compression ratio as a percentage, and encoding speed as well as decoding speed in samples per microsecond.
+The full dataset, [undefn], contains 50+ entries.
+Each entry specifies additional metadata (e.g., genre) as well as original file size, number of digital samples, compression bit rate, compressed file size, original encoding algorithm, and encoding time as well as speed for encoding and decoding for each of the tested audio codecs.
+Codec Setting Compression Ratio Encoding Speed Decoding Speed
 Lossless FLAC level 6 38.74 31.92 92.47
@@
 Table 4: Average audio-codec compression as percent ratio and encoding as well as decoding speeds in samples per microsecond. Best performance in each category has been marked.
+Table 4 presents average values for compression ratio, encoding speed and decoding speed of our dataset.
 FLAC, the only lossless compression scheme tested, achieves an average encoding speed of 32 samples per microsecond, i.e.
@@
 4.2 Perceptual Quality
-Figure 2 visualizes the spectra of all tested audio codecs versus the uncompressed digital audio file of the song Iron Man by Black Sabbath.
 Figure 2: Spectra for all tested audio codecs (exported from Audacity).
-X-axis denotes the frequency range up to 24 kHz.
-The y-axis shows the loudness range from 0 dB to -150 dB.
-The color-coded graphs refer to uncompressed (black), FLAC level 6 (red), MP3 CBR 320 kbps (yellow), MP3 CBR 128 kbps (green), AAC CBR 256 kbps (orange), AAC VBR level 5 (teal), and Vorbis VBR level 7 (pink).
-The graphs represent encoded audio signal outputs for the respective audio codec.
-Note the spectrum for the uncompressed audio signal is completely aligned with the spectrum of FLAC level 6, showing that FLAC encoded audio exhibits the exact same loudness as the uncompressed audio signal across the entire frequency range.
-The graph of FLAC level 6 aligns perfectly with the spectrum of uncompressed audio, showing that the loudness of FLAC encoded audio exactly matches the loudness of the uncompressed audio across the entire spectrum.
+X-axis denotes frequency range up to 24 kHz.
+Y-axis shows loudness range from 0 dB to -150 dB.
+Color-coded graphs refer to uncompressed (black), FLAC level 6 (red), MP3 CBR 320 kbps (yellow), MP3 CBR 128 kbps (green), AAC CBR 256 kbps (orange), AAC VBR level 5 (teal), and Vorbis VBR level 7 (pink).
+Graphs represent encoded audio-signal outputs for the respective audio codec.
+Note the spectrum for the uncompressed audio signal is completely aligned with that of FLAC level 6, meaning both exhibit the exact same loudness across the entire frequency range.
+Spectra of both MP3, AAC CBR 256 kbps, and Vorbis encoded audio diverge from uncompressed audio spectrum, revealing the impact of lossy compression schemes on the overall loudness.
+Figure 2 visualizes the spectra of all tested audio codecs versus the uncompressed digital audio of Iron Man by Black Sabbath.
+The graph for FLAC level 6 aligns perfectly with the spectrum of the uncompressed audio signal, showing that the loudness of FLAC encoded audio exactly matches the loudness of the uncompressed audio across the entire spectrum.
 On the other hand, spectra for all audio encoded with lossy compression techniques deviate from the spectrum of the original audio signal.
 The same pattern of reduction in loudness can be seen for the interval 15 kHz to above 20 kHz.
-MP3 320 kbps and MP3 128 kbps elevate the loudness of the entire frequency range before cutoff at 18 kHz and 15 kHz, respectively.
+MP3 320 kbps and MP3 128 kbps elevate the loudness of the entire frequency range before cutoff occurs at 18 kHz and 15 kHz, respectively.
 AAC CBR 256 kbps loudness reduction happens much earlier in the frequency range, around 7 kHz.
-To our surprise, AAC VBR level 5 loudness closely matches with the spectra for FLAC and the uncompressed audio very well with only a slight loudness reduction at the high-frequency range.
+To our surprise, AAC VBR level 5 loudness closely matches with the spectra for FLAC and uncompressed audio with only a slight loudness reduction at the high-frequency range.
 Vorbis level 7 also exhibits loudness reduction at high frequencies.
-However, unlike the other tested lossy compression techniques, it manifests a smooth decline in loudness.
-[SpekWebsite] was used to further investigate and visualize the loudness change in encoded audio compared to uncompressed audio.
-Figure 3 shows that all audio encoded with lossy compression schemes exhibit very quiet loudness in the range -100 dB to -105 dB, marked dark blue in figure 3a, which is further reduced to near silence at -120 dB, marked black in figures 3c to 3g, for high frequencies above 15 kHz.
+However, unlike the other tested lossy compression techniques it manifests a smooth decline in loudness.
 Figure 3: Audible loudness comparison for the track Iron Man by Black Sabbath using different encoding schemes.
@@
 Spectrograms indicate loudness in dB.
-MP3 320 kbps retains some portion of very quiet loudness until 20 kHz, while AAC VBR level 5 retains some very quiet loudness. Vorbis level 7 loudness reduction is not as flat as other lossy codecs.
-Figure 4 shows example sound-field visualizations for stereo images and sound stages capturing the 3:00 – 3:30 min interval of Iron Man by Black Sabbath, when Tony Iommi performs the guitar solo.
+[undefad] was used to further investigate and visualize the loudness change in encoded audio compared to uncompressed audio.
+Figure 3 shows that all audio encoded with lossy compression schemes exhibit very quiet loudness in the range -100 dB to -105 dB, marked dark blue in figure 3a, which is further reduced to near silence at -120 dB, marked black in figures 3c to 3g, for high frequencies above 15 kHz.
+MP3 at 320 kbps maintains much of the original structure but exhibits slight attenuation in the upper frequencies, while MP3 at 128 kbps shows a more pronounced high-frequency cutoff and reduced spectral detail.
+AAC and Vorbis demonstrate similar high-frequency constraints and reshape frequency-dependent energy in the mid-to-lower frequency.
+The redistribution of quantization noise is part of AAC’s and Vorbis’ perceptual encoding strategy to compress audio content while making sure the audio quality is perceptually masked [undefe].
 Figure 4: Sound-field visualizations for the track Iron Man by Black Sabbath compressed using different audio encoding schemes.
-LABEL:sub@fig:evaluation:iron+man+sf:a shows the sound-field visualization from 3:00 – 3:30 min for the uncompressed signal and LABEL:sub@fig:evaluation:iron+man+sf:b – LABEL:sub@fig:evaluation:iron+man+sf:g show the sound-field visualization for the same time interval encoded with FLAC, MP3s, AACs, and Vorbis, respectively.
-The patterns of dots should lean to the right, capturing the stereo image of the electric guitar’s sound in the recording.
-Compared to the uncompressed audio signal’s stereo image, the FLAC encoded audio signal was able to accurately capture the stereo image while audio signals from other lossy encoders show some inaccuracies.
-Listening to the song, the guitar solo sounds slightly to the right, so the stereo image of this session should lean to the right of the center as well.
-All the sound-field visualizations in figures 4b to 4g have their clusters lean to the correct direction.
+LABEL:sub@fig:evaluation:iron+man+sf:a shows the sound-field visualization from 3:00 – 3:30 min for the uncompressed audio signal and LABEL:sub@fig:evaluation:iron+man+sf:b – LABEL:sub@fig:evaluation:iron+man+sf:g show the sound-field visualization for the same time interval encoded with FLAC, MP3, AAC, and Vorbis, respectively.
+Dot patterns should lean to the right, capturing the stereo image of the electric guitar’s sound in the recording.
+Compared to the uncompressed audio signal’s stereo image, FLAC encoded audio was able to accurately capture the stereo image while other lossy encoders exhibit some inaccuracies.
+Figure 4 shows example sound-field visualizations for stereo images and sound stages capturing the 3:00 – 3:30 min interval of Iron Man by Black Sabbath when Tony Iommi performs the guitar solo.
+Listening to the uncompressed audio the guitar solo appears slightly to the right, which leads to the expectation that the stereo image of this session should lean to the right of the center as well.
+As expected the sound field for the FLAC encoded audio in figure 4b is nearly indistinguishable from that of the uncompressed audio.
+All the sound-field visualizations in figures 4b to 4g have their clusters lean in the expected direction.
 However, the clusters in figures 4c to 4g, representing audio encoded with lossy compression techniques, significantly deviate from the uncompressed sound field in figure 4a.
 The sound-field clusters are too sparse for figures 4c to 4d using MP3 encoding, while the clustering for AAC CBR 256 kbps in figure 4e is too condensed in the center.
-However, both AAC VBR level 5 (cf.
+However, sound-field clusters for both AAC VBR level 5 (cf.
 figure 4f) and Vorbis level 7 (cf.
-figure 4g) sound-field clusters are very similar to the uncompressed audio sound field in figure 4a.
-Finally, advanced PEAQ ODG scores and totalNMR were calculated based on GstPEAQ [Holters:GSTPEAQ:2015] and [peaqbGitHub], respectively, to quantify the impact of compressing digital audio signals on perceptual quality.
-Table 5 provides advanced PEAQ ODG scores and totalNMR values for all tested samples in table 3.
-Advanced PEAQ ODG score range is 0 to -4 table 2 and the lower totalNMR the better.
-Track Codec Setting Adv. ODG totalNMR
-Dreams Fleetwood Mac FLAC level 6 0.2 -131.08
-MP3 320 kbps -0.1 -0.77
-MP3 128 kbps -0.1 -1.07
-AAC 256 kbps -0.1 0.15
-AAC level 5 -0.1 0.15
-Vorbis level 7 -0.1 -16.75
-La mer II. Jeux de Vagues Claude Debussy FLAC level 6 0.2 -121.33
-MP3 320 kbps -2.0 -2.04
-MP3 128 kbps -2.4 -2.4
-AAC 256 kbps -1.4 -1.03
-AAC level 5 -1.4 -1.03
-Vorbis level 7 -0.1 -17.56
-Foreigner Pallbearer FLAC level 6 0.2 -126.49
-MP3 320 kbps -0.6 -3.04
-MP3 128 kbps -0.2 -3.33
-AAC 256 kbps -0.2 -2.35
-AAC level 5 -0.2 -2.34
-Vorbis level 7 -0.2 -17.28
-September Earth, Wind & Fire FLAC level 6 0.2 -135.91
-MP3 320 kbps -0.2 -1.14
-MP3 128 kbps -3.5 -1.43
-AAC 256 kbps -0.2 -0.46
-AAC level 5 -0.2 -0.46
-Vorbis level 7 -0.1 -16.74
-Table 5: Advanced PEAQ ODG score and totalNMR for all tested samples in table 3.
-Table 6 presents the average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs at their most common quality settings.
-Codec Setting Adv. ODG totalNMR
-FLAC level 6 0.2 -130.11
-MP3 CBR 320 kbps -0.2 -1.34
-MP3 CBR 128 kbps -0.4 -1.72
-AAC CBR 256 kbps -0.4 -0.08
-AAC VBR level 5 -0.4 -0.08
-Vorbis VBR level 7 -0.1 -16.74
-Table 6: Average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs.
-While all audio codecs score near 0 on advanced PEAQ ODG, FLAC’s average totalNMR is significantly lower than the average totalNMR of all lossy codecs, with Vorbis comes in second.
-All audio codecs score near 0 on average advanced PEAQ ODG score.
-However, FLAC’s average totalNMR is the lowest.
-Vorbis’s average totalNMR is second lowest but there is a significant gap between FLAC and Vorbis.
-The remaining average totalNMR are -1 to 0.
+figure 4g) are very similar to the uncompressed audio sound field in figure 4a.
+Track Codec Setting BPEAQ ODG 2f Score APEAQ ODG BPEAQ totalNMR
+Dreams Fleetwood Mac FLAC level 6 0.2 100 0.2 -131.08
+MP3 320 kbps -3.8 32 -0.1 -0.77
+MP3 128 kbps -3.8 32 -0.1 -1.07
+AAC 256 kbps -3.9 29 -0.1 0.15
+AAC level 5 -3.9 29 -0.1 0.15
+Vorbis level 7 0.0 100 -0.1 -16.75
+La mer II. Jeux de Vagues Claude Debussy FLAC level 6 0.2 100 0.2 -121.33
+MP3 320 kbps -3.1 60 -2.0 -2.04
+MP3 128 kbps -3.5 62 -2.4 -2.40
+AAC 256 kbps -3.6 56 -1.4 -1.03
+AAC level 5 -3.8 65 -1.4 -1.03
+Vorbis level 7 0.0 100 -0.1 -17.56
+Foreigner Pallbearer FLAC level 6 0.2 100 0.2 -126.49
+MP3 320 kbps -3.2 71 -0.6 -3.04
+MP3 128 kbps -3.6 71 -0.2 -3.33
+AAC 256 kbps -3.6 63 -0.2 -2.35
+AAC level 5 -3.5 63 -0.2 -2.34
+Vorbis level 7 0.0 100 -0.2 -17.28
+September Earth, Wind & Fire FLAC level 6 0.2 100 0.2 -135.91
+MP3 320 kbps -3.5 29 -0.2 -1.14
+MP3 128 kbps -3.6 30 -3.5 -1.43
+AAC 256 kbps -3.7 27 -0.2 -0.46
+AAC level 5 -3.8 27 -0.2 -0.46
+Vorbis level 7 0.0 100 -0.1 -16.74
+Table 5: Scores for basic PEAQ ODG, 2f model, and advanced PEAQ ODG as well as basic PEAQ totalNMR values for all tested audio tracks in table 3.
+Our dataset was then tested using multiple PEAQ experiments and overall results for a variety test tracks are shown in table 5.
+As expected, lossless FLAC achieves perfect scores under all metrics.
+This means FLAC’s perceptual quality is near-identical to the original audio’s perceptual quality.
+Both MP3 320 kbps and MP3 128 kbps, along with AAC 256 kbps and AAC VBR level 5, score near -4 on BPEAQ, indicating very perceptible impairment (cf.
+table 2).
+The 2f score of both MP3 codecs as well as both AAC codecs range from 27 (poor) to 71 (fair) indicating that, if these audio codecs were tested in MUSHRA, trained listeners would find these four audio codecs somewhat decent on some tracks and below average on other tracks.
+The observed differences in 2f scores may partly reflect variations in the spectral and temporal characteristics of the source material.
+A pop song like Fleetwood Mac’s Dreams as well as material from the disco genre, e.g., Earth, Wind & Fire’s September, contain dense rhythmic instrumentation including prominent percussion and sustained mid–to-high frequency energy, which increase temporal modulation complexity and may amplify codec-induced modulation distortion.
+In contrast, a classical performance such as Debussy’s La mer exhibits more structured orchestral textures with comparatively smoother spectral evolution.
+Rock performances like Pallbearer’s Foreigner contain heavily distorted guitar signals whose broadband harmonic content can raise masking thresholds and partially conceal additional codec artifacts.
+These genre dependent signal properties may contribute to the better 2f scores observed for the latter tracks.
+Using APEAQ, most encoded tracks show good ODG score of 0 and -1.
+The only exceptions are Debussy’s La mer, scoring at -2 for MP3 320 kbps and MP3 128 kbps, and Earth, Wind & Fire’s September, scoring worst at -3.5 for MP3 128 kbps, indicating perceptible impairment (cf.
+table 2).
+This indicates that the compression impairment in MP3 and AAC audio codecs may not be as severe as previously shown by BPEAQ on some tracks.
+However, audio impairment of Earth, Wind & Fire’s September encoded with MP3 128 kbps is rated in both BPEAQ and APEAQ as quite perceptible.
+totalNMR values for both MP3 audio-codec settings as well as both AAC audio-codec settings range from -3 to 0 meaning the noise to mask ratio is rather high and most noise from these encoded test tracks will be audible to listeners.
+The most interesting case in table 5 is Vorbis VBR level 7.
+It is the only lossy audio codec that scores 0 on both BPEAQ ODG and APEAQ ODG as well as 100 on 2f score.
+This strongly indicates Vorbis’ audio impairment is imperceptible to listeners and the audio codec’s perceptual quality is very close to lossless FLAC perceptual quality.
+Vorbis’ totalNMR value is around -17, which means its noise level is less audible than the remaining tested audio codecs’ noise level.
+Codec Setting BPEAQ ODG 2f Score APEAQ ODG BPEAQ totalNMR
+FLAC level 6 0.2 100 0.2 -130.11
+MP3 CBR 320 kbps -3.6 36 -0.2 -1.34
+MP3 CBR 128 kbps -3.7 36 -0.4 -1.72
+AAC CBR 256 kbps -3.7 32 -0.4 -0.08
+AAC VBR level 5 -3.8 37 -0.4 -0.08
+Vorbis VBR level 7 0.0 99 -0.1 -16.74
+Table 6: Average scores for basic PEAQ ODG, 2f model, and advanced PEAQ ODG as well as PEAQ totalNMR values for all tested audio codecs (cf.
+table 6).
+Best performance in each category has been marked.
+Table 6 presents the average BPEAQ ODG scores, average 2f scores, average APEAQ ODG scores as well as average BPEAQ totalNMRs for all tested audio codecs using the quality settings defined in table 6.
+BPEAQ ODG for FLAC is 0 and for Vorbis 0.2 on average, translating to perceptual quality (nearly) the same as the original audio signal.
+The remaining audio codecs’ BPEAQ ODG average close to -4, which means theses audio codecs’ impairment is classified as very annoying (cf.
+table 2).
+These partial results are supported as well by the average 2f-model scores, which indicate that FLAC and Vorbis would score 100 on the MUSHRA listening test while the remaining codecs perform at much lower values in the 30s.
+The average totalNMR also shows that noise in all lossy audio codecs is very noticeable except for Vorbis.
+The interesting case here is the APEAQ ODG score, which shows much better values for all audio codecs at near 0.
+This indicates that APEAQ regards all tested audio codecs impairment as imperceptible, which contradicts with all the BPEAQ ODG scores, 2f-model scores, and totalNMR values.
 4.3 Machine-Learning Based Codec
 In recent years, a trend emerged to apply machine-learning techniques into remastering and compressing audio.
-One proposed audio compression model employs a residual vector quantized generative adversarial network (RVQGAN) [Kumar:HFACIRVQGAN:2023].
-RVQGAN works by encoding the audio signal into a lower-dimensional presentation followed by converting this presentation into discrete values.
-We employed RVQGAN’s prototype implementation from its [DACRVQGANGithubRepo] and encoded the track Iron Man by Black Sabbath.
-Table 7 presents RVQGAN’s average compression ratio, average encoding speed and average decoding speed together with traditional audio codecs for comparison.
-Codec Setting Comp. Ratio Enc. Speed Dec. Speed
+One proposed audio compression model employs a residual vector quantized generative adversarial network (RVQGAN) [undefw].
+RVQGAN works by encoding the audio signal into a lower-dimensional presentation followed by conversion into discrete values.
+We employed RVQGAN’s prototype implementation from its [undefg] and encoded the Black Sabbath’s Iron Man with it.
+Codec Setting Compression Ratio Encoding Speed Decoding Speed
 Lossless FLAC level 6 38.74 31.92 92.47
@@
 Table 7: Average audio-codec compression as percent ratio and encoding as well as decoding speeds in samples per microsecond.
-Values for lossless and lossy encoding are the same as in table 4 plus the results for RVQGAN.
+Values for lossless and lossy encoding are the same as in table 4 with an additional result row for RVQGAN.
 Best performance in each category has been marked.
-In its current state RVQGAN does improve on the compression ratio by reducing storage size to almost 1% of the original size.
+Table 7 presents RVQGAN’s average compression ratio, average encoding speed, and average decoding speed in comparison with traditional audio codecs previously evaluated.
+In its current state, RVQGAN does improve on the compression ratio by reducing storage size to almost 1% of the original size.
 However, encoding and decoding speed are nearly one to two magnitudes slower than any of the traditional audio codecs.
-Figures 5 to 7 illustrate the spectrum, spectrogram, and sound-field visualization of an RVQGAN encoded audio signal compared to the original uncompressed audio signal.
 Figure 5: RVQGAN’s spectrum (red) together with the uncompressed signals spectrum (blue) (exported from Audacity).
@@
 It is apparent that the pattern in the RVQGAN encoded audio signal output is more smeared and lacks the accuracy of the guitar’s position when compared to the uncompressed audio’s sound field.
+Figures 5 to 7 illustrate the spectrum, spectrogram, and sound-field visualization of an RVQGAN encoded audio signal compared to the original uncompressed audio signal.
+The spectrogram in figure 6 shows a cutoff in loudness at around 21 kHz.
+The spectrum in figure 5 shows the RVQGAN tested track exhibits first a slight loudness increase in the lower frequency range, which gradually decreases form around 10 kHz to -100 dB at 20 kHz and beyond.
+figure 7 shows the encoded audio-signal output pattern more smeared and less accurate compared to the uncompressed track’s original pattern.
+Overall, these visualizations indicate that RVQGAN encoded audio signals may share similar perceptual quality problems with lossy encoded audio.
+Codec Setting BPEAQ ODG 2f Score APEAQ ODG BPEAQ totalNMR
+FLAC level 6 0.2 100 0.2 -130.11
+MP3 CBR 320 kbps -3.6 36 -0.2 -1.34
+MP3 CBR 128 kbps -3.7 36 -0.4 -1.72
+AAC CBR 256 kbps -3.7 32 -0.4 -0.08
+AAC VBR level 5 -3.8 37 -0.4 -0.08
+Vorbis VBR level 7 0.0 99 -0.1 -16.74
+RVQGAN -2.7 37 -3.7 -0.69
+Table 8: Average scores for basic PEAQ ODG, 2f model, and advanced PEAQ ODG as well as PEAQ totalNMR values for all tested audio codecs (cf.
+table 6) plus average scores for RVQGAN.
+Best performance in each category has been marked.
 Despite achieving an impressive compression ratio of around 98% of the uncompressed file size, perceptual quality of RVQGAN encoded audio is questionable.
 Currently, RVQGAN exhibits worse perceptual quality than all traditional audio codecs.
-RVQGAN’s average ODG score and average totalNMR in table 8 supports this evaluation.
-Codec Setting Adv. ODG totalNMR
-FLAC level 6 0.2 -130.11
-MP3 CBR 320 kbps -0.2 -1.34
-MP3 CBR 128 kbps -0.4 -1.72
-AAC CBR 256 kbps -0.4 -0.08
-AAC VBR level 5 -0.4 -0.08
-Vorbis VBR level 7 -0.1 -16.74
-RVQGAN -3.7 -0.69
-Table 8: Average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs. This is table 6 with the additional advanced PEAQ ODG score and totalNMR for RVQGAN.
+RVQGAN’s average evaluation scores in table 8 supports this evaluation.
 5 Discussion
@@
 In our audio-codec performance tests (cf.
-table 4), FLAC exhibited the lowest compression ratio of all tested audio codecs, meaning that lossless audio files will take up more storage space than files encoded using a lossy compression technique.
-On the other hand, FLAC also achieved the highest encoding speed.
-These two average values for FLAC are expected because lossless compression techniques preserve most data in an audio file while only removing redundancies and repeated patterns.
-All lossy audio codecs have similar compression ratio, with MP3 (128 kbps) achieving the highest compression ratio.
-Both MP3 quality settings exhibit lower encoding speeds than other lossy audio codecs.
-On the other hand, they show significantly higher decoding speeds than other lossy audio codecs with MP3 at 128 kbps being the fastest to decode.
+table 4), FLAC exhibited the lowest compression ratio of all tested audio codecs meaning that lossless audio files will take up more storage space than files encoded using a lossy compression technique.
+On the other hand, FLAC achieved the highest encoding speed.
+These two average values for FLAC are expected because lossless compression techniques preserve most of the information in an audio signal while only compressing redundancies and repeated patterns.
+All lossy audio codecs have similar compression ratios with MP3 128 kbps achieving the best results.
+Both MP3 quality settings, i.e.
+128 kbps and 320 kbps, exhibit lower encoding speeds than other lossy audio codecs.
+On the other hand, these codecs show significantly higher decoding speeds than other lossy audio codecs with MP3 128 kbps exhibiting the fastest decoding speed.
 It can be argued that MP3 is a strong contender for the best lossy audio codec due to its high decoding speeds, making it the ideal choice for audio playback on low-power devices.
+This is because MP3 has lower computational complexity compared to modern codecs like AAC or Vorbis, which is due to its older more simple perceptual encoding model [undefc].
+However, MP3’s encoding speed is constrained by its two-nested-iteration-loop design [undefc], which makes MP3’s encoding speed fall behind when comparing to AAC and Vorbis newer and more sophisticated encoding techniques.
 Among lossy audio encoders, Vorbis exhibits above average compression ratios, the fastest encoding speeds, and decent decoding speeds.
-In our audio-codec performance evaluation, Vorbis is the runner up for lossy audio-codecs, providing decent compression ratios and fastest lossy encoding speeds.
+In our audio-codec performance evaluation, Vorbis is the runner up for lossy audio codecs, providing good compression ratios and fastest lossy encoding speeds.
 5.2 Perceptual Quality
-Our perceptual quality evaluation shows that lossy compression algorithms sacrifice high-frequency aural details to reduce storage size.
-Audio encoding using psycho-acoustic principles determines these region are mostly inaudible to the average listener and the audio data for these regions can simply be removed.
-The differing graphs in figure 2 for all lossy audio codecs lies in how each codec handles loudness reduction of the high-frequency region.
-Loudness reduction or removal of high frequencies reduces the harmonic richness, the details and clarity, and the spatial depth [Truax:THFAE:2020] that allow digital music to exhibit high fidelity audio.
-[Vickers:TLWBSR:2010], mainly reporting on the loudness war in music production, also shows that the average human listener can perceive aural quality very differently depending on the loudness alteration in a music recording.
-Additionally, loudness reduction and high-frequency removals affect the stereo image of the encoded audio reducing their aural accuracy.
-The APEAQ ODG scores and totalNMR values we collected from all tested audio compression codecs (cf.
-table 6) support our findings.
-On average, all traditional audio codecs - FLAC, MP3, AAC, and Vorbis - score near 0 on APEAQ ODG.
-This means that APEAQ evaluates all of these traditional audio codecs impairment level as imperceptible (cf.
+Our perceptual quality evaluation shows that, while lossy compression algorithms still try to preserve perceptual audio quality, they sacrifice high-frequency aural details to reduce storage size, thus reducing the quality of the audio signal.
+Because audio encoding using psycho-acoustic principles determines regions that are mostly inaudible to the average listener, audio data for these regions are simply removed.
+The spectra for all lossy audio codecs (cf.
+figure 2) clearly show how each lossy audio codec handles loudness reduction of high-frequency regions.
+Loudness reduction and removal of high frequencies reduces the harmonic richness, the details and clarity, and the spatial depth [undefah] that allow digital music to exhibit high-fidelity audio qualities.
+[undefai], mainly reporting on the loudness war in music production, also shows that the average human listener is able to perceive aural quality very different depending on the loudness alteration in a music recording.
+Additionally, loudness reduction and high-frequency removals negatively affect the stereo image of the encoded audio signal reducing aural accuracy (cf.
+figure 1c).
+We employ multiple complementary PEAQ predictors, including basic PEAQ, advanced PEAQ, basic PEAQ’s totalNMR values, and the 2f model.
+Using a combination of not just ODG scores but also MOV scores allows to capture distinct psychoacoustic dimensions.
+Convergence across these metrics (cf.
+table 6) strengthens confidence that observed audio-codec artifacts align with listener perceptible phenomena, providing a reproducible computational approximation rather than a substitute for subjective evaluation.
+The ODG scores and the MOV scores collected from all tested audio compression codecs (cf.
+table 6) support these findings.
+On average, all traditional audio codecs, i.e.
+FLAC, MP3, AAC, and Vorbis, score near 0 on advanced PEAQ ODG.
+This means that advanced PEAQ classifies all of these traditional audio codecs’ impairment level as imperceptible (cf.
 table 2).
-However, the noise level of lossy codecs is higher than noise level of lossless codecs.
-In particular, noise levels for MP3 and AAC are right at the threshold to their mask levels, meaning their noise levels are potentially perceptible.
-Remarkably, Vorbis is the only lossy audio-compression codec where the noise level is masked reasonably well, making it the winner of this category.
-While most of our assessment was conducted via a system of codec performance measurements, visualizations, and PEAQ tests, we have to address the importance of listening tests.
-PEAQ were built under [ITU:BS.1387-2:2023] to predict how an average human perceive audio quality [ITU:BS.1387-2:2023].
-Trained listeners, under [ITU:BS.1534-3:2015] on conducting MUSHRA tests, need to meet certain requirements to be qualified as assessors [ITU:BS.1534-3:2015].
-Thus, they can provide important and practical feedback about how humans perceive aural quality, which validate the result of the objective evaluation models, hence validate our study.
+However, the noise level of lossy audio codecs is higher than the noise level of lossless audio codecs.
+In particular, noise levels for MP3 and AAC are right at the threshold of their masking levels, meaning their noise levels are potentially perceptible.
+Remarkably, Vorbis is the only lossy audio-compression codec where the noise level is masked reasonably well, making it the winner in this category.
+The 2f-model scores support this as well with Vorbis’ achieving an average score of 99 while other lossy encoders’ average 2f-model scores are in the 30s.
+The notable discrepancy between basic PEAQ ODG scores and advanced PEAQ ODG scores observed for encoded audio with MP3 and AAC (cf.
+table 6) warrants further discussion.
+While basic PEAQ yields an average ODG score of approximately -3.7 for both codecs, suggesting very annoying impairment, advanced PEAQ in contrast produces scores near 0, suggesting imperceptible impairment.
+This gap can be attributed to the difference in perceptual modeling depth between the two models.
+Basic PEAQ’s MOVs, which rely primarily on noise-to-mask ratio and loudness-based metrics [undefaf], can be rather rigid when evaluating audio artifacts above basic PEAQ’s masking threshold, even if not be perceptible for a human listener.
+Advanced PEAQ, by contrast, incorporates MOVs capturing modulation masking, harmonic coherence, and asymmetric distortion weighting [undefaf], which together better account for the structured, partially-masked nature of audio encoding artifacts.
+The above findings suggests that basic PEAQ’s ODG score should be interpreted with caution when evaluating modern perceptual codecs.
+The magnitude of the gap between basic ODG scores to advanced ODG scores in itself may serve as an indicator of how well a codec’s artifact profile aligns with a given perceptual model’s assumptions.
+Our finding of a gap between basic PEAQ ODG scores and advanced PEAQ ODG scores also agrees with work by [undefj], which motivated conducting additional evaluation with the 2f model and basic PEAQ’s totalNMR MOV.
+Together these four metrics allow a comprehensive and objective comparison of an audio codec’s perceptual quality characteristics.
+While most of our assessment was conducted via a system of audio codec performance measurements, visualizations, and various PEAQ-related tests, we have to address the importance of subjective listening tests.
+PEAQ variants were built under [undefr] to predict how an average human listener perceives audio quality [undefr].
+Trained listeners, under [undefs] conducting MUSHRA tests, are obligated to meet certain (minimum) requirements to qualify as assessors [undefs].
+Thus, trained listeners may be able to provide additional feedback about how humans perceive aural quality, which would be useful to validate the results of the objective evaluation models and, hence, validate and improve our study and approach.
 5.3 AI-based Audio Codecs
-Regarding AI-based compression technique, it is known that there were several neural-network models that were proposed.
-These models include an end-to-end approach [Rim:DNNETE:2021], lightweight transformer approach [Défossez:HFNAC:2022], and a novel audio codec called APCodec [Ai:APCodec:2024].
-We chose the RVQGAN model by [Kumar:HFACIRVQGAN:2023] because the authors claimed their model outperforms competing traditional audio compression algorithms.
-Our tests indicate RVQGAN indeed exhibits the highest compression ratio compared with traditional audio codecs (cf.
+Our tests indicate RVQGAN [undefw] indeed exhibits the highest compression ratio compared with traditional audio codecs (cf.
 table 7).
@@
 However, RVQGAN’s encoding and decoding speeds are very low compared to traditional audio codecs, performing one to two magnitudes slower.
-Additionally, in its current version, RVQGAN’s output is a unique format descript-audio-codec (.dac) that is not playback-supported by any audio player application at the time of our evaluation tests.
-While the RVQGAN codec achieved the highest average compression ratio in our audio-codec performance tests, Our testing shows that RVQGAN’s perceptual quality is worse than traditional audio codecs.
-RVQGAN’s average advanced ODG score is -3.7 which is close to very annoying (cf.
-table 2), and RVQGAN’s average totalNMR indicates that its noise level is quite perceptible (cv.
+Additionally, in its current version RVQGAN’s output is a unique file format, descript-audio-codec (.dac) [undefg], which is not playback-supported by any audio player application at the time of our evaluation tests.
+While the RVQGAN codec achieved the highest average compression ratio in our audio-codec performance tests, our testing shows that RVQGAN’s perceptual quality is at the worse end of traditional audio codecs’ perceptual quality.
+RVQGAN’s average advanced ODG score is -3.7, which is close to very annoying (cf.
+table 2), and RVQGAN’s average totalNMR value indicates that its noise level is quite perceptible (cf.
 table 8).
-The AI-based approach to developing audio codecs shows promise in advancing efficiency but will require further refinements to provide acceptable hi-fi listening experiences to end users.
+AI-based approaches to developing or discovering audio codecs shows promise in advancing efficiency but will require further refinements to provide acceptable hi-fi listening experiences to end users.
 6 Conclusions
-We evaluated several of the most common audio-compression codecs providing insights in how audio compression techniques affect sonic perceptual quality of digital music.
-Our tests indicate that lossy audio-compression codecs do provide better compression performance than lossless audio compression codecs.
+We evaluated several of the most common audio-compression codecs providing insights in how audio-compression techniques affect sonic perceptual quality of digital music.
+As expected, our tests indicate that lossy audio-compression codecs do provide better compression performance than lossless audio-compression codecs.
 This stems mainly from the fact that lossy audio-compression algorithms sacrifice aural perceptual quality for higher compression speeds and lower final storage size.
 However, the Vorbis audio codec emerged as the exception to this rule.
-It exhibits acceptable compression efficiency when compared to its competitors while at the same time achieving perceptual quality nearly imperceptibly different compared to respective uncompressed digital audio signal.
-We also evaluated RVQGAN, an AI-based audio compression scheme.
-While RVQGAN’s codec performance in speed and size was impressive, its perceptual quality is severely lacking.
-We are sure to see more advancement in this field, especially improving the perceptual quality of future AI-driven audio compression codec developments.
-Our evaluation strengthens our argument that, when choosing a digital audio compression codec, focus should not only be on the audio compression performance.
+It exhibits acceptable compression efficiency when compared to its competitors while at the same time achieving perceptual quality nearly imperceptibly distinguishable compared to the respective uncompressed digital-audio signal.
+Finally, if storage size is not a concern, lossless audio-compression codecs such as FLAC will be provider better aural perceptual quality in reproducing the original audio signal.
+We also evaluated RVQGAN, a machine-learning based audio compression scheme.
+While RVQGAN’s codec performance in encoding speed and final storage size was impressive, its aural perceptual quality is severely lacking.
+We are sure to see more advancement in this field, especially improving the perceptual quality of future AI-driven audio-compression codec developments.
+Our evaluation strengthens our argument that, when choosing a digital-audio compression codec, not only audio-compression performance should be a focus.
 Users should also consider the perceptual quality of the encoded audio signal to ensure the best possible listening experience.
-We compliment Vorbis as an excellent lossy audio-compression codec also able to maintain high-fidelity perceptual quality.
+We compliment Vorbis as an excellent lossy audio-compression codec, which is also able to maintain high-fidelity perceptual quality.
+The demonstrated consistency between objective evaluation metrics, i.e.
+vasic PEAQ, advanced PEAQ, and the 2f model, and established MUSHRA behavior provides further evidence that objective perceptual modeling can serve as a robust foundation for audio-codec evaluation without reliance on extensive listening panels.
 6.1 Future Work
-Expanding the database of audio codec performance measurements can determine the consistency in compression ratio, encoding speed, and decoding speed of all audio compression codec.
-We welcome readers to add performance measurements to our database https://github.com/samcate/fmedia-Audio-Codec-Comparison-Database.
-With the recent trend of high-fidelity music streaming services like [Spotify], perceptual quality evaluation should be conducted on streaming audio.
-Streaming services utilize a wide range of audio-compression schemes, which can provide interesting insights.
-Future research should also focus on enhancing perceptual quality in AI-driven compression methods.
-By integrating psycho-acoustic principles into the compression neural network, it may be possible to achieve more accurate perceptual quality, providing low-storage size high-fidelity digital music to users.
+Expanding our database of audio-codec performance measurements can determine the consistency in compression ratio, encoding speed, and decoding speed of all audio-compression codecs.
+We welcome additional performance-measurement submissions to our database https://github.com/samcate/fmedia-Audio-Codec-Comparison-Database by the community.
+With the recent trend of increased use of high-fidelity music streaming services like [undefae], perceptual quality evaluation should be extended to streaming audio as well.
+Streaming services utilize a wide range of audio-compression schemes, which can provide interesting insights, especially when employing objective scoring.
+Future research should also focus on enhancing the perceptual quality of AI-driven compression methods.
+By integrating psycho-acoustic principles into a compression neural network it may be possible to achieve more accurate perceptual quality in addition to providing low-storage size high-fidelity digital music to users.